# Constraints on Compact Dark Matter Population from Micro-lensing Effect of Gravitational Wave for the third-generation gravitational Wave Detector

### v1

@@ IV Results
For the population hyperparameters of compact DM, we fix the minimum mass $M_{\min}=10\leavevmode\nobreak\ M_{\odot}$ and maximum mass $M_{\max}=10^{3}\leavevmode\nobreak\ M_{\odot}$ in the mass function of compact DM, and we adopt uniform priors for the power-law index $\gamma\in\mathcal{U}[-1,3]$ and abundance of compact DM $f_{\rm CO}\in\mathcal{U}[-4,-1]$ in Eq. (16), respectively. Then we incorporate the above-mentioned 8 inferred lens masses from micro-lensing GW events into the EMCEE [49] with the posterior Eq. (16) to estimate population hyperparameters $\Phi\in[\gamma,\leavevmode\nobreak\ f_{\rm CO}]$. Our results are shown in Figure. 3. We obtain that the best-fit values and $68\%$ confidence levels for the hyperparameters $\gamma,\leavevmode\nobreak\ f_{\rm CO}]$ are $\gamma=0.41^{+0.42}_{-0.37}$, $\log_{10}(f_{\rm CO})=-2.76^{+0.16}_{-0.17}$, respectively. As shown by the blue data point in Figure 3, which indicate the truth values for hyperparameter $\Phi\in[\gamma=1,\leavevmode\nobreak\ \log_{10}(f_{\rm CO})=-2.5]$ in our simulations, our posterior estimates demonstrate consistency with these input truth values within $2\sigma$ confidence levels.
V Conclusion
@@
In this paper, we present a comprehensive methodology for: (1) simulating GW events mciro-lensed by compact DM, (2) extracting information about both the GW sources and lens for each micro-lensing events, and (3) constraining the population properties of compact DM from these events. Specifically, by combining the future third-generation ground-based GW detector ET with point mass model, we calculate the posterior distribution of GW sources and lens parameters for the detection of 8 micro-lensing GW events among $10^{4}$ observed BBH mergers. Then, based on the hierarchical Bayesian inference framework, we demonstrate that for a population characterized by a power-law mass function, these 8 micro-lensing GW events would constrain the power-law index to $\gamma=0.41^{+0.42}_{-0.37}$ and compact DM abundance to $\log_{10}(f_{\rm CO})=-2.76^{+0.16}_{-0.17}$ at $68\%$ confidence levels as shown in Figure 3. Our analysis demonstrates agreement between the derived results and the true input values from simulations within $2\sigma$ confidence levels. This discrepancy is expected, as our theoretical predictions suggest that more than 10 micro-lensing GW events should be detectable given the true parameter values. However, only 8 micro-lensing GW events were detected, which explains why the derived constraints on both $f_{\rm CO}$ and $\gamma$ are lower than the true input values. The implementation of our methodology will play a pivotal role in future observations of micro-lensing GWs: it enables not only the precise extraction of population characteristics of compact DM, but also provides a pathway to probe their fundamental nature—whether they stem from primordial black holes, exotic compact objects, or other non-particle dark matter candidates.

### v2

@@ V Results
For the population hyperparameters of compact DM, we fix the minimum mass $M_{\min}=10\penalty 10000\ M_{\odot}$ and maximum mass $M_{\max}=10^{3}\penalty 10000\ M_{\odot}$ in the mass function of compact DM, and we adopt uniform priors for the power-law index $\gamma\in\mathcal{U}[-1,3]$ and abundance of compact DM $f_{\rm CO}\in\mathcal{U}[-4,-1]$ in Eq. (35), respectively. Then we incorporate the above-mentioned 8 inferred lens masses from micro-lensing GW events into the EMCEE [61] with the posterior in Eq. (35) to estimate population hyperparameters $\Phi\in[\gamma,\penalty 10000\ f_{\rm CO}]$. Our results are shown in Figure. 3. We obtain that the best-fit values and $68\%$ confidence levels for the hyperparameters $\gamma,\penalty 10000\ f_{\rm CO}]$ are $\gamma=0.41^{+0.42}_{-0.37}$, $\log_{10}(f_{\rm CO})=-2.76^{+0.16}_{-0.17}$, respectively. As shown by the blue data point in Figure 3, which indicate the truth values for hyperparameter $\Phi\in[\gamma=1,\penalty 10000\ \log_{10}(f_{\rm CO})=-2.5]$ in our simulations, our posterior estimates demonstrate consistency with these input truth values within $2\sigma$ confidence levels.
VI Conclusion
@@
In this paper, we present a comprehensive methodology for: (1) simulating GW events mciro-lensed by compact DM, (2) extracting information about both the GW sources and lens for each micro-lensing events, and (3) constraining the population properties of compact DM from these events. Specifically, by combining the future third-generation ground-based GW detector ET with point mass model, we calculate the posterior distribution of GW sources and lens parameters for the detection of 8 micro-lensing GW events among $10^{4}$ observed BBH mergers. Then, based on the hierarchical Bayesian inference framework, we demonstrate that for a population characterized by a power-law mass function, these 8 micro-lensing GW events would constrain the power-law index to $\gamma=0.41^{+0.42}_{-0.37}$ and compact DM abundance to $\log_{10}(f_{\rm CO})=-2.76^{+0.16}_{-0.17}$ at $68\%$ confidence levels as shown in Figure 3. Our analysis demonstrates agreement between the derived results and the true input values from simulations within $2\sigma$ confidence levels. This discrepancy is expected, as our theoretical predictions suggest that more than 10 micro-lensing GW events should be detectable given the true parameter values. However, only 8 micro-lensing GW events were detected, which explains why the derived constraints on both $f_{\rm CO}$ and $\gamma$ are lower than the true input values. The implementation of our methodology will play a pivotal role in future observations of micro-lensing GWs: it enables not only the precise extraction of population characteristics of compact DM, but also provides a pathway to probe their fundamental nature, whether originating from primordial black holes, exotic compact objects, or other non-particle dark matter candidates.
Although our method has established a preliminary framework for constraining dark matter using future gravitational wave data, it still needs to account for additional factors to better approximate the real physical process, such as
  • While our study focuses exclusively on the point-mass model, compact DM in the universe may exist in other extended halo structures, such as “dressed PBH” [62, 63, 64, 65, 66]. If an observed signal is a true micro-lensing waveform, using an incorrect template model will produce a biased result. Therefore, following the detection of a micro-lensing GW event, it is crucial to accurately identify the specific lens model involved [35, 51].
  • Conversely, the second case arises when an observed waveform is not lensed but is spuriously identified as a micro-lensing signal. A prime example is the analysis of GW from a binary in an elliptical orbit using templates designed for quasi-circular orbits. In such instances, the micro-lensing analysis may mistakenly interpret the elliptical orbit signature as evidence of lensing. Consequently, additional methods are required to break this degeneracy with the parameters of lens, for example, employing the Kramers-Kronig (KK) relation [67].
  • To date, no GW event has shown unambiguous signs of micro-lensing. Future detections will require more complete waveform models to robustly interpret any deviations and to assess the micro-lensing contribution from known stellar populations [68, 69, 70, 71, 72]. Ultimately, determining whether we can reliably identify dark matter substructures, distinct from the effects of known baryonic matter, remains an open challenge.